Amiga-FAQ

Frequently asked questions and answers

29 November 1993

Jochen Wiedmann

Copyright \begin{matrix}c \\ ◯ \end{matrix} Jochen Wiedmann

Am Eisteich 9

72555 Metzingen (Germany)

Tel. 07123 / 14881

Internet: wiedmann@mailserv.zdv.uni-tuebingen.de

This document lists some frequently asked questions and tries to give answers. Is intention is to help new users and to reduce the amount of news that most experienced users don’t like to read anymore.

Permission is granted to make and distribute verbatim and modified copies of this document following the terms of the “GNU General Public License” provided the copyright notice and this permission notice are preserved on all copies.

The author gives absolutely no warranty that the answers given here are correct. Many of them were contributed by other users and I cannot even make marginal checks. If you think that something should be changed, please tell me. Suggestions, contributions, new answers, critics, flames (oh, how I like this ‘nil:’ :-) are rather welcome.

Please notice that there are many questions that aren’t answered yet, even more: Whole sections that remain empty! I feel that I cannot give satisfying answers. So it’s your turn: Fill the gaps and tell me what i should include into this document!

1 Hardware

This chapter contains questions concerning the Amiga-Hardware.

1.1 What are the 68EC020 and the 68EC030?

Motorola, the company producing the 680x0 family offers crippled versions of their processors. They are a little bit cheaper than the originals, that’s why Commodore decided to build the 68EC020 into the A1200 and the 68EC030 into the A4000/030.

The difference between the 68020 and the 68EC020 is that the latter can address just 16Mb of memory. That’s why the A1200 cannot have more that 10 Mb RAM. In most cases you will not notice the difference.

This is not the case for the 68EC030: Many owners will notice that the 68030 has a MMU and the 68EC030 doesn’t have. There are some important programs depending on a MMU, for example Enforcer (a debugging utility), GigaMem (a program to emulate virtual memory) or all current Unix versions (see Section 6.1 [Unix], page 16). Other Amigas than the A4000 and the A3000 need an additional processor card to run these.

1.2 What’s a FPU?

The first 680x0 processors (upto 68030) could process integers only. Floating point operations had to be emulated by the software. A FPU is a chip (or part of a chip) that can process floating point operations, a mathematical coprocessor.

One separates three FPU types on the Amiga: The 68881, 68882 and the 68040’s internal FPU. The 68882 is up to 1.5 times faster than the 68881, because it is splitted in two parts: A conversion unit (the FPU’s are using an 80 bit format internally) and the arithmetic unit. The 68040’s internal FPU adds a pipeline, but misses the trigonometric instructions of the others. These are still emulated by the software, 68040.library for example.

Special programs (Raytracing, DTP, Mathematics, TeX) are offered in a special coprocessor version which are up to 50 times faster than the original versions.

Michael Kaiser (kaiser@ira.uka.de)

2 The Operating System

This chapter handles questions concerning the operating system, Kickstart as well as the Workbench.

2.1 Can I use another Kickstart than the builtin?

First let’s drop some words on the Kickstart’s Copyright: This belongs to Commodore, you must not use Kickstarts, without the right to do it! Especially it isn’t allowed to make an image of anyone else’s Kickstart and run this on your own Amiga. (I even doubt that it is allowed to do this on your own Amigas, if you have more than one.)

But of course it is possible and allowed for some people, developers for example. There are two different ways, a hardware solution and a software solution. The former is to buy a card which can hold two or more Kickstart ROMS and allows to select between when the System is booting.

The software solution needs a program (softkicker) and an image of the ROM. The softkicker allocates RAM, loads the ROM image into the allocated memory and reboots. Of course you have less RAM after the Reboot: 256Kb when running Kickstart 1.2 or 1.3 and 512Kb for Kickstart 2.0 or higher.

There are different Softkickers, some of them needing a MMU (see Section 1.1 [MMU], page 1). A nearly perfect solution seems to be kick13 (Aminet, directory ‘util/misc’) because it has the ROM image included. (With Commodores agreement!) For newer Kickstarts I recommend SKick 3.43 (Aminet, directory ‘os20/util’) because it doesn’t need a MMU and supports many different Kickstarts. It is rather easy to create the ROM image using the following program:

         #include <stdio.h>
         #define kickorig 0xf80000 /*  0xfc0000 for Kick 1.2 und 1.3   */
         #define kicklen  0x080000 /*  0x040000 for Kick 1.2 und 1.3   */
         void main(int argc, char*argv[])
           { FILE *fh;
     if ((fh = fopen("kickstart.file", "w"))  !=  NULL)
       { result = fwrite(kickorig, kicklen, 1, fh);
       }
     fclose(fh);
           }

3 Programming

This chapter handles problems arising for programmers only.

3.1 What documentation do I need as an Amiga programmer?

The best information available are the RKM’s (ROM Kernel Manuals), 3rd edition, by Commodore, published by Addison-Wesley:

The Amiga ROM Kernel Manual: Libraries The Amiga ROM Kernel Manual: Devices The Amiga ROM Kernel Manual: Includes and Autodocs The Amiga Hardware Reference Manual The Amiga User Interface Style Guide

Especially the RKM: Libraries is a must. The RKM: Includes and Autodocs isn’t that much worth: Better get the same stuff on disk instead to have it online. See Section 3.3 [Includes], page 4.

AmigaDOS isn’t included in these books. The Autodocs give much information, but to go deeper you probably need

The AmigaDOS Manual, 3rd Edition, by Commodore

published by Bantam Books.

Another good choice is The Amiga Guru Book from Ralph Babel. The book starts with a survey on different aspects of programming the Amiga. (About 250 pages) Useful for beginners are the sections on the Amiga’s data types (not the 3.0-DataTypes), the Includes and the amiga.lib. But even experienced programmers will find useful things here that are missing in the RKM’s. But the largest part are about 500 pages on AmigaDOS and, as I think, the most important, because AmigaDOS is the worst officially documented part of the OS. The book is rather concise and hence not as easy to read as the RKM’s, but I recommend it as an addition and instead of the AmigaDOS manual. (However, no replace for the Libraries and Devices, which aren’t handled here.) Unfortunately the book has no ISBN and is currently available in Germany only. But there are some mail order companies which offer it for about 50$ and which accept credit cards:

Buchhaus Gonski, Neumarkt 18a, 50667 Koeln, Germany
Phone: 0221/2090972,    Fax: 0221/2090959
Buchhandlung Bouvier, Am Hof 32, 53113 Bonn, Germany
Phone: 0228/7290169     Fax: 0228/7290178
Hirsch & Wolf OHG, Mittelstrasse 33, 56564 Neuwied, Germany
Phone: 02631/83990      Fax: 02631/839931
(Eurocard/Mastercard/VISA)
Mainhattan-Data, Schoenbornring 14, 63263 Neu-Isenburg, Germany
Phone: 06102/5881       Fax: 06102/51525
(Eurocard/Mastercard/VISA/American Express)
DTM-Computersysteme, Dreiherrenstein 6a, 65207 Wiesbaden, Germany
Phone: 06127/4064       Fax: 06127/66276
(Eurocard/Mastercard)
Unlimited GmbH, Kehrstrasse 23, 65207 Wiesbaden, Germany
Phone: 06127/66555      Fax: 06127/66636

3.2 What is CATS?

This is a department at Commodore West Chester which was formerly named Commodore Amiga

Technical Support and was later renamed Commodore Application and Technical Support.

These are people that work independently of Engineering, but close together with them, and try to help developers outside of Commodore to create nice Amiga applications, software or hardware. To achieve this, CATS has gathered a lot of informations and tools, on floppy, CD, or paper. Much of this material is also available to the general public.^{1} But don’t mix this up with some sort of Hotline for everyone!

For Americans the address to get this material is

         Commodore Business Machines, Inc.
         Department C
         1200 Wilson Drive
         West Chester, PA 19380,

for all Europeans it’s a company in Germany:

         Fa. Hirsch & Wolf
         Mittelstr. 33
         56564 Neuwied
         Tel. 02631/83990

Dr. Peter Kittel, peterk@cbmger.de.so.commodore.com

3.3 Where do I get the Amiga includes?

The only legal way to get the includes and autodocs (and you should get them, they are very useful!) is to buy the NDUK (Native developers update kit), which is offered by CATS. They cost about 25$ and this seems to me to be a fair price. See Section 3.2 [CATS], page 4.

Unfortunately there is only the NDUK for Kickstart 2.0. People say that 3.0 will come but nobody knows when. But there are two legal possibilities, to get at least the 3.0-Includes (without the AutoDocs, sigh!):

1. Some Fish-CD’s (those with the installed software) contain them. See Section 8.4.2 [Fish CD], page 23.

2. Commercial C compilers and the registered Dice version have the 3.0-Includes. Dice is a good and cheap offer for about 50$. See Section 3.5 [Compilers], page 5.

Finally you could become a developer. But that is expensive. See Section 3.4 [Developer], page 5.

1 Which means: For Non-Developers.

3.4 How do I become a developer?

You need the ADSP (Amiga Developer Support Program) documents. To get this write a letter to your local Commodore branch asking for these documents. Everything else should be explained there. The german address is

Commodore
Lyoner Strasse 38
60528 Frankfurt

There are three different developer versions:

Registered developers seem to me to have no special advantages except getting a magazine called AmigaMail. Especially they don’t receive the new Includes & AutoDocs and beta releases of Kickstart and Workbench. Being a registered developer costs about 70$ per year.

Certified developers don’t have these disadvantages. They have to sign the so called NDA (Non Disclosure Agreement) and so aren’t allowed to tell other people what they know. Certified developers pay about 280$ per year.

Commercial developers pay about 550$ per year.

A tip is to build a group of users and become a developer group, so the costs are reduced.

3.5 What compilers (assemblers) are there?

There is a lot of programming languages on the Amiga, commercial as well as freely distributable. I will enumerate only those that I know or which seem it worth to me otherwise.

Assembler All C-compilers have an Assembler included. Freely distributable are A68K and PhxAss (directory ‘dev/asm’ on Aminet or Fish disks 521 and 906)

C C++ Freely distributable C-compilers are gcc (which has its own directory ‘dev/gcc’ on

Aminet) and the evaluation version of Dice (for example per FTP from ftp.uni-

paderborn.de, directory ‘/news/comp.binaries.amiga/volume91/languages’ or on

Fish disk 491). Dice is Shareware, however, but registering will cost only 50$. The advantage of gcc is that you find gcc versions all over the world and on all computer systems. Another advantage is that C++ is included into gcc! But it is slow and needs 4Mb of RAM or more. Dice is very fast. Both compilers seem to be reliable. Commercial C compilers are Aztec-C and SAS-C. I cannot recommend Aztec-C, because the compiler doesn’t seem to get further development. It should be remarked that the commercial compilers have especially wonderful debugging utilities (Source level debuggers!) that the others are missing. Dice will become commercial soon too and probably will have debugging utilities then. SAS has announced a new release which will have C++ included.

Comeau C++ is a cross compiler creating C source. That wouldn’t be a problem, SAS will do the same. But Comeau C++ doesn’t have a C compiler included. You need SAS-C, Aztec-C or Dice additionally. But it supports the newest standards. And like gcc it runs on many platforms. Maxxon C++ is offered in Germany. I cannot say anything on it. Both compilers are commercial. Comeaus address is:

         Comeau computing
         91-34, 120th Street
         Richmond Hill, NY, 11418-3214
         USA
         EMail: Greg Comeau, comeau@bix.com

Forth JForth is said to be an excellent Amiga port of Forth. Among its advantages are object oriented extensions, full Amiga interface and an application generator. It is available from:

         Delta Research
         P.O. Box 151051
         San Rafael, CA   94915-1051
         Phone: (415) 453-4320
         EMail: Phil Burk, phil@ntg.com
        Mike Haas, haas@starnine.com

Fortran (Sigh! Still people who need it :-<) Freely distributable are BCF (Fish disk 470) and f2c, a Fortran to C converter (Aminet, directory ‘/dev/misc’). A commercial compiler is offered from ABSoft. All these are Fortran 77 compilers, I don’t know any Fortran 90 compiler on the Amiga.

Lisp Freely distributable Lisp interpreters are XLisp (Fish disk 181) and OakLisp (Fish disk 530).

Prolog  ‘/dev/lang/UNSWProlog.lha’ and ‘dev/lang/sbp3_1e.lha’ on Aminet as well as

‘SBProlog’ on Fish disk 141 and ‘SBProlog’ on Fish disk 145 are freely distributable Prolog interpreters.

Modula-2 M2Amiga is offered in Europe, Benchmark Modula-2 in the U.S. Both are said to be very good, have a powerful souurce-level-debugger, a large library. Especially M2Amiga has great support by a german user-group (AMOK) which for example offers own PD disks.

M2Amiga is offered by:

                   A+L AG
                   Daderiz 61
                   2540 Grenchen
                   Schweiz
                   Tel.: +41/65/52 03-11
                   Fax:              -79

Benchmark Modula-2 is available from:

                   Armadillo Computing
                   5225 Marymount Drive
                   Austin, Texas 78723
                   USA
                   Phone/FAX: 512/926-0360.
                   EMail: Jim Olinger, jolinger@bix.com

Oberon Oberon-2 AmigaOberon is offered by A+L too. It is integrated into a full developers environ-

ment and has a large library of modules. Library linker and source level debugger are available. The AMOK user goup supports AmiOberon as well as M2Amiga.

Pascal There is a PD-compiler called PCQ (Directory ‘dev/lang’ on Aminet or Fish disk 511). It doesn’t support all of Pascal and major features are missing. P2C, a pascal to C converter is on disk 341. (Aminet: ‘/dev/misc/p2c120.lha’) Additionally there are two commercial compilers called HiSoft Pascal and KickPascal. HiSoft Pascal and P2C claim to be compatible to Turbo Pascal up to 5.0. HiSoft has a source level debugger included.

3.6 Those never working Esc sequences!

Many printers come with a manual that explains which ESC sequence causes which action on the printer. But there happen weird things when you try to send these sequences to your printer, either it does nothing, or it does something completely different. There is a reason, the Amiga printer drivers. These drivers are made in a way that they only understand a certain set of ANSI Esc sequences, not the special ones defined (differently) by the various printer manufacturers. The purpose is that every application on the Amiga just uses this one standard set of control sequences and this way doesn’t need to know which printer is actually connected. The printer driver then translates these standard sequences into the special sequences a certain printer understands. A list of the available ANSI Esc sequences is found in the current Workbench manuals (or older AmigaDOS manuals). Now if you want to issue a control sequence to the printer that’s not available as an ANSI command, you have two possibilities to achieve this:

1. Bypass the printer driver (that would unsuccessfully try to interpret the sequence) and send your output only during this sequence to ‘PAR:’ (or ‘SER:’, respectively). For this you have to close and open printer output channels very often which is rather tedious, and you have to know where (‘PAR:’ or ‘SER:’) your printer is connected.

2. Use a special ANSI sequence, made exactly for this case:

             `Esc[<n>"<x>'

where ‘<n>’ is the decimally typed number of bytes in the string ‘<x>’, which actually contains your special printer sequence. This ANSI sequence tells the printer driver to not interpret or translate the next ‘<n>’ bytes.

But both methods have one big disadvantage when used in an application program: You lose the printer independency! If you stick to ANSI sequen- ces, you can output to any printer on earth, as long as there is an Amiga printer driver for it. If you start to use special control sequences, your program will be tied to this single printer model and will not be usefull for any other (or you would have to provide some dozen new printer drivers for your application).

Dr. Peter Kittel, peterk@cbmger.de.so.commodore.com

3.7 Is it possible to use AmigaBasic on the A1200?

We heard conflicting reports about AmigaBasic on the A1200: While I said that you can work sufficiently with it, others said that this isn’t possible, as AmigaBasic crashes on the slightest little error. I couldn’t reproduce this.

Now I can. It depends on the setting in the Sound Prefs editor. When you activate a sound there, this conflicts with sound that AmigaBasic tries to produce by hand and obviously not quite the correct way itself.

Easy solution: To work with AmigaBasic on the A1200, just

1. Switch off sound output in the Sound prefs editor.

2. On the A4000 (as well as on an A1200 with Fast Mem expansion) you additionally need to run NoFastMem.

3. Better avoid SUBs and use conventional GOSUBs instead, then the compatibility with newer processors will be higher.

Dr. Peter Kittel, peterk@cbmger.de.so.commodore.co

4 Applications

This chapter offers informations about major Applications.

4.1 Text Editors

Text Editors are programs allowing to enter and edit unformatted text. Generally, this means text that is meant to be manipulated by machine, rather than human. Programmers use these to enter the text for compilers. Since UNIX machines don’t typically have word processors, most text processing starts with a text editor, then is filtered through a page layout system (Section 4.4 [LaTeX], page 10, for example) to produce attractive paper results.

Commercial Products CygnusEd Professional and TurboText seem to be the main contenders in the professional realm. The Fred Fish disks contain dozens of other shareware text editors. A demo version of TurboText is on Fish disk 445. A very old demo of CygnusEd is on Fish disk 95 (testament to its lasting-power). In the following some freely distributable editors will be discussed.

Emacs Gnu Emacs (the "G" is not silent) comes from Unix and is probably the king of editors – it’s huge (about 1 Megabyte), feature-packed (it does windows and even contains a game!) and extensible (if you know lisp you can write new emacs functions and bind them to any key combination). On the other hand, it may be too huge, its featureladenness is imposing, and its extensibility often means you can’t use someone else’s emacs configuration. Source: Aminet (directory ‘util/gnu’).

Vi The leaner, less configurable, non-extensible cousin to gnu emacs is vi (pronounced "vee eye"). Unix people like vi especially because you find it on any Unix machine. What you choose is personal preference, and will mark you for life. Vim is a good vi for the amiga, and is on Fish disk 591 or in the ‘util/gnu’ directory of Aminet.

DME Many Amiga programmers like DME. It’s fast, fully configurable; menus may be created and any key may be mapped. It’s much easier to learn DME than than Emacs or Vi. There are three different versions: AmokEd, DME and XDME. It’s a matter of opinion which you prefer. (Oberon and Modula programmers like AmokEd because it’s written in Oberon and supports AmiOberon error messages, C programmers like DME or XDME.) Sources: Aminet (directory ‘util/edit’), Fish disk 776 (XDME) and 749 (AmokEd), AMOK 90.

4.2 What word processors are there?

A word processor is the typical application for writing notes, letters or reports on a computer. Unless you prepare newsletters on a weekly basis, your word procesor is probably your workhorse program. Thus choosing one you are comfortable with determines how comfortable you are with your computer. Word processors can offer a variety of features, and many can approach the sophistication required for Desktop Publishing (see Section 4.3 [DTP], page 10) but no one uses them for programming, for which text editors are more suited.

One distinguishes between Wysiwyg programs (What you see is what you get) and page layout languages. Wysiwyg programs should be fast, comfortable and easy to use. Most people prefer them. The alternative is an approach that works similar to compilers. You feed text files to a a program that produces the layout which may be previewed on screen or printed. LaTeX takes this approach. See Section 4.4 [TeX], page 10. Lout is another such system which seems smaller, easier to learn and has full documentation included, but it is nonstandard. Lout produces Postscript output. (I don’t know if this is an advantage or disadvantage. ;-) See Section 4.5 [Postscript], page 11. Both programs are freely distributable.

There are a lot of wysiwyg programs, but only commercial products: Final Copy II, Wordworth, Word Perfect, AmiWrite, Beckertext II, Maxon Word and many others. I don’t dare to recommend any. ALl I can say is: Give yourself time to make a selection.

4.3 Desktop Publishing

These programs offer features lacking in word processors, usually tailored to flexible arrangement of text, but often don’t provide all of the text manipulation that a good word processor provides. The best desktop publishing programs strive to provide the features of both, just as the best word processors strive to provide the features of desktop publishing programs. Microsoft Word (Mac, PC) is a good example of a word processing program that offers many page layout feature. Framemaker (UNIX, Macintosh, DOS, etc.) is an example of a desktop publishing system that offers most needed word processing functions. As yet, no Amiga program has bridged the gap, though the main word processors are coming close. (On the other hand, even many sophisticated programs don’t support typesetting mathematics, tables, producing bibliographies, indexes, or cross-references. The page layout languages do, and programs like Frame are improving their support of such features.) Unless you need to prepare fancy newsletters or promotional literature, a word processing program is probably enough. See Section 4.2 [Word Processors], page 9.

There are not yet any freely distributable wysiwyg desktop publishing systems. Commercial products are ProPage and PageStream. They have been playing leapfrog for the past few years. It appears that PageStream 3.0 is about to leap ahead. A more detailed description of these products and their differences is welcome. Both programs’ list prices are $299. Student discounts are available (approx 40% discount.)

4.4 What is TeX and where can I get it?

TEX is a very powerful wordprocessing system. It can display mathematical formulas or complex tables as well as function graphs, creates indices, contents and many other things. Its greatest advantage is that it is freely distributable (TeX, not the previewers and the printer drivers!) and that you find TeX all over the world on every computer family. Its greatest disadvantage is that it isn’t very handy (works similar to a compiler) and it isn’t wysiwyg. But many people like it. (BTW: This document is written using TeX. 8-) See Section 4.2 [Word Processors], page 9.

There are two major implementations on the Amiga. The first one, Amiga-TEX, from Thomas Rockicki and Radical Eye software is commercial. It is said to be excellent and his owners seem to be very satisfied. But it costs at least 200$.

I recommend PasTEX, a freely distributable version. People seem to have problems installing PasTeX, especially the font loading and generation (It’s a quite complex program.) but I did not here anyone upset once it was installed. (A friend with knowledge of TeX helps immensely.) A few words should be said what you need:

• 5 disks containing the TEX-compiler itself

• 2 disks containing MetaFont

• Nothing more

Many people ask for fonts. They are included in the MetaFont-package and can get compiled by you. All you need to do is setting up your TEX-system in the right way which is described in the documentation. Please note that the PasTeX disks are compressed using the program Zoom. (see Section 7.2 [Endings], page 17) Sources: FTP at ftp.uni-passau.de, directory

‘/pub/amiga/tex/PasTeX1.3’.

4.5 Are there any Postscript interpreters?

PostScript is a programming language designed to be used to describe printing on pages. Apple helped make PostScript popular by selling printers with built in PostScript interpreters. Many programs have evolved to produce PostScript programs as their output, making PostScript the lingua franca of printing. Until recently, in order to print a PostScript file, you had to have a relatively expensive laser printer. The development that changed this was the software PostScript interpreter. These programs allow your computer to interpret PostScript programs, and produce the matrix of dots to send to your normal graphics printer.

One of the benefits of PostScript is that it is resolution independent. What this means is that it can support the highest resolution of your device – and that you can reasonably preview PostScript on a low resolution screen.

There are two free PostScript interpreters for the Amiga. Post and Ghostscript. Post comes as an Amiga shared library along with front ends for previewing to the screen and printing. This structure allows others to write programs that can show PostScript images on screen. In fact, AmigaTeX uses Post’s library to support incorporation of PostScript into documents. Ghostscript similarly comes in two programs, but not as a shared library. Ghostscript is the rendering engine, and Ghostview is the front end. Sources: Aminet (directorys ‘text/print’ and ‘text/dtp’), Fish disk 669

5 How about Graphics?

Graphics is one of the major strengths of the Amiga. Why don’t we have more answers here? :-(

5.1 What are chunky and planar displays?

Simply put, the terms chunky and planar (short for bitplanar) refer to different ways of storing graphics information in a computer’s memory. They are rather easy to understand, as far as things go, but incredibly difficult to explain:

Computer images are arranged as a grid of pixels, each of which can be thought of as a number representing the color number of the pixel, sort of like a paint-by-numbers scheme. For example, here’s a simplified example image, in four colors:

        00302132

The Amiga stores this image in a bitplane mode. That is, it is represented by several planes of bits (binary digits, 1s or 0s). This is a four-color image, so each color number could be represented by two bits. Therefore there are two bitplanes:

        00100110    Here's bitplane 0
        00101011    And here's bitplane 1
        --------    Now, let's add them up, binary style:
        00302132

Which is the final image. If the image was in two dimensions, it would truly be composed of bit planes. However, I’d need three dimensions to show multiple bitplanes overlayed, and therefore for simplicity we’re working in one dimension (which is all we need).

Now, there’s another way of storing this image. How about if we localize the bit data in little chunks?

        00 00 11 00 01 10 11 01 = 00302132

This is the principle of the chunky pixel mode.

Both methods of image storage are perfectly logical, and no one can say that one is better than the other. However, there are certain technical aspects which cause certain advantages and disadvantages.

First, if you’ve seen colored text scroll on your Amiga, you know there is a bit of "flicker" that arises. Specifically, what happens is that while the text is scrolling, its color temporarily changes to something completely different. What’s happening is that the computer’s moving several bitplanes of data while the raster (monitor electron gun) is sweeping across the screen. What that means is that, if the raster catches the data while it’s being moved, you can end up with some bitplanes being moved and some not. What if we filled bitplane 1 in the example above with 0s? Instantly all the 3s become 1s, and the 2s become 0s! This is what causes "flicker" when certain colors are scrolled. By contrast, if a chunky pixel display is caught while scrolling, all we see is a partially-scrolled image; the colors are preserved (since their units are the small ones).

That’s a disadvantage to planar pixels, but what about chunky pixels? Well, recall that a computer organizes information in terms of 8 bit bytes. These groups are static; you cannot decide to all of a sudden organize data in terms of three bytes or something! Therefore, when using chunky pixels, things get complicated if we decide to use a nonconvenient number of bits per pixel. In practice, the 8-bit (256-color) mode, and 24-bit (16 million color) modes are the most common candidates for chunky pixel displays.

Finally, certain effects can be accomplished with the different systems. Bitplanar mode is particularly useful for things like shadows (where an extra bitplane is set with 1s instead of 0s), and chunky mode is great for perspective and "mapping" (since the data for each pixel is localized in a single "chunk"). The latter advantage makes chunky pixel mode really great for games, and is what made Wolfenstein 3-D possible.

We all know that Amigas use the bitplane system for storing images. However, the Macintosh and PC(VGA) both use chunky pixel modes. While we can optimize our RAM usage with "bizarre" modes like 8- and 128-color, they gain the advantages of non-flicker scrolling, and the programming simplicity of just writing a byte where you want the pixel to go.

The difference between the two modes becomes problematic in things like emulation. EMPLANT has a "chunky to planar" routine which it uses to convert a Macintosh display into an Amiga one. "Chunky to planar" routines are also useful for getting chunky-inclined things to run on Amigas (see TMAPDemo, rotdemo). On a side note, there was some confusion as to what EMPLANT used the MMU for with regard to chunky to planar. The MMU itself is incapable of performing the algorithm for the conversion; rather, it is used to detect what portions of the display memory are updated from the Mac side, and therefore the processor is saved from having to perform the chunky to planar conversion for the entire display.

I sincerely hope that helped clear up most of the mystery concerning the terms "Chunky" and "Planar"!

(Joseph Luk, jluc@eis.calstate.edu)

5.2 What is doublebuffering?

Don’t be ashamed if you haven’t a clue as to what this is. It’s rather simple, really. Imagine you wanted to animate something by drawing out each frame. Now imagine you had only one piece of paper. Even if you drew pretty fast, it would be difficult to see what the animation looks like because you’d have to erase the frame you just drew, in order to draw the next one!

What if you had two pieces of paper? Why, then, you could see the difference between two adjacent frames. First you’d draw your first frame on the first piece of paper, then you’d draw the second frame on the other piece. Then you’d erase the picture on the first piece of paper, and draw in the third frame. Then you’d erase the picture on the second piece of paper, and draw in the fourth frame, and so forth.

This is the principle of doublebuffering. The computer shows you the picture it just drew, then turns around and draws the next one. It then swaps the picture in front of you with the one it just drew, and repeats the process. The result is a smooth animation, because you never need to see the computer draw; all you see is each finished product.

(Joseph Luk, jluc@eis.calstate.edu)

5.3 What monitors will work on my Amiga 1200/4000?

Monitors can be classified after the horizontal scan frequency they require. TVs, as well as C=’s 1084 monitor, need frequencies around 15 kHz. VGA/SVGA need approx. 30 kHz. Multisync monitors can take many frequencies.

In short: You can use any monitor you want with an A1200. BUT:

− If you use a regular VGA/SVGA monitor, you can only use a few display modes (like DblPAL, DblNTSC and/or Productivity). I.e. (320|640) x (256|512|1024) for DblPAL. This is great for Workbench and all "serious" utilities (DTP etc), but don’t expect any games to work... they don’t use your preferences, just take over the machine and assume a 15 kHz monitor. Also, you cannot utilize the "Early Startup Control" screen (you know, disable cache, and that stuff), which also requires a 15 kHz monitor. Furthermore, VGA monitors don’t have speakers. And the VGA-type modes don’t support Genlocks. But for a lot of "serious" work, a VGA monitor is quite adequate.

− You already know what happens with a 15 kHz monitor; the flickering in Interlace mode. A small tip: Try to use NTSC instead of PAL. This increases the refresh rate from 25 Hz to 30 Hz, at the expense of lower vertical resolution (482 lines maximum). The NTSC and PAL modes aren’t as bad as many people think. If your monitor has a lot of phosphorous (long afterglow), PAL Laced can be quite OK, and it gives you a resolution of 1448x566 in SuperHiRes. That’s the highest resolution currently supported on AGA Amigas, in any display mode.

− A Multisync gives you the best of both worlds. The new 1940 and 1942 monitors from C= are quite OK, although rather cumbersome to use... The h/v size and offset must be set manually each time you switch display mode.^{2}

For a VGA/SVGA or Multisync monitor, you’d need a little shiny box which gives you the standard 15-pin "D" connector. It costs around $15.

There are other alternatives... like the "AmiVGA" box ($50 or so), which I think is a cheapo version of the Flicker Fixer. (But it’s really a shame to use this on an AGA Amiga.)

Also, you can get a VGA monitor, and hook up your TV to the composite or RF port on the Amiga - one monitor for games, one for serious stuff.

(Per Espen Hagen, per.e.hagen@ffi.no)

5.4 How do I switch between PAL and NTSC?

PAL and NTSC are two different video standards, the former being European, and the latter being American. PAL has a slightly taller screen (256 lines non-interlaced, non-overscanned) as opposed to NTSC (200 lines), so if you see the bottom portion of a program’s screen getting cut off on your American machine, chances are the program was written for PAL, and is running on your shorter NTSC screen. PAL and NTSC differences are somewhat less important to European users; since their machines default to PAL, running an NTSC program is no more than a minor annoyance having the screen only appear in the top portion of the display.

Therefore, for us NTSC folks, switching into PAL mode becomes important to avoid loss of some picture on Euro Demos, etc.

First, the most common misconception about switching between PAL and NTSC is that you need a Multiscan or special monitor for such a purpose. Not so! Just about any monitor can handle the minor signal difference between PAL and NTSC (50Hz vertical refresh versus 60Hz). The 108x, 19xx, and 20xx series, and even most TVs, can display both PAL and NTSC. Sometimes it is necessary to perform minor tweaking of vertical hold and/or v. size on your monitor to achieve full display, though this is trivial.

What IS needed to switch between PAL and NTSC in software, is a "Fatter" (1MB) Agnus or better. If your system has more than 512K of CHIP RAM (use the avail command to find out) you have this chip. If you do not, ou can still construct a hardware switch (see below).

The following directions assume you’re an NTSC user who wants to switch into PAL mode, but the procedure for going from PAL to NTSC is much the same.

2 A patch is available on Aminet for Kickstart 3.0 to do this without manual actions in the file

   ‘os30/util/Monitor30Patch.lha’.

The most common use of switching into PAL is for self-booting games, demos, etc. The best way to accomplish this on pre-3.0 systems is to use Chris Hames’ Degrader program (most recent version: 1.30). Once you have procured this program, the switch is as simple as selecting "50Hz", and "50Hz System", then installing the program’s ROMTag (little program that runs at boot-time) by pressing the "Survive Reset(s)" button.

AmigaDOS 3.0 added a PAL/NTSC switch feature to its Boot Menu, and all you need to do in order to access this is hold down both mouse buttons as the computer boots. Select Display Options, Display Mode/PAL, and then Boot. If the program still fails to go into PAL mode (Zool is one I’ve found), you may wish to get Degrader anyway and try that.

Software PAL/NTSC switching is easy and painless. Some people have gone even further by installing hardware PAL/NTSC switches. This results in a system which even the most nasty hardware-banging programs can’t bring to its original configuration (if it has been switched). The procedure is simple – most Amigas have jumpers already – but beyond the scope of this document.

Therefore, if your screen is too short or cut off, seek the different screen mode. You’ll "see new horizons", literally!

(Joseph Luk, jluc@eis.calstate.edu)

6 Emulators

What? The Amiga isn’t good enough? You really want it to be another machine? Well, look here…

6.1 Can I run Unix on my Amiga?

Actually there are three Unix versions on the Amiga. All of them need at least 68030 (see Section 1.1 [MMU], page 1), probably a 68040 in the near future and seem to have problems with many hard-drive-controllers. You should have a good look into the documentation before installing it. Unix needs much resources, say at least 10Mb RAM and a 150Mb Unix-partition on the hard-drive.

1. Commodore offered a System V Unix in the past. It contained TCP/IP, X11 and other software and seemed to make a good job, but it was expensive and after all Commodore has dropped developing it. Commodore-Unix needs a streamer because it is distributed on tapes.

2. A Linux port is prepared. But actually there is not very much than the kernel. Specialists might like to use it, but it cannot be recommended for now. Linux is freely distributable.

    Sources: ftp.uni-paderborn.de, directory ‘/pub/amiga/linux’ or ftp.uni-erlangen.de,
    directory ‘/pub/LINUX/MIRROR.tsx-11/680x0’.

3. NetBSD is freely distributable too. Like Linux it isn’t ready for now, but it seems to make big steps. Most GNU software is said to run, especially emacs and gcc. I think it’s worth to have a look on it. Sources: ftp.uni-paderborn.de, directory ‘/pub/amiga/NetBSD’ or

    ftp.uni-erlangen.de, Directory ‘/pub/amiga/unix/NetBSD’.

6.2 Is it possible to use the Amiga as X11 terminal?

Yes, it is. There are two different packages available:

GfxBase offers a commercial version which is said to be excellent, but expensive. The distribution includes different window managers andclients. A demo version is on Aminet.

(‘gfx/x11/GfxBase-X11-Demo.lha’)

DaggeX is freely distributable and probably not finished yet. (It calls itself version 0.22.) Source:

Aminet, ‘gfx/misc/localDaggeX0_22.lha’ and ‘gfx/x11/twm_930531.lha’.

7 Miscellaneous

This last chapter contains some questions that don’t fit in the chapters above.

7.1 Is there any unix version of LhA?

See Section 7.2 [Endings], page 17.

7.2 What are files ending with ...?

Most endings on FTP sites or Fish disks tell you that the file is compressed and/or is an archive containing more than one file. Some programs even archive whole disks. Frequently found endings and programs to handle the related files are:

.lha .lzh Compressed archives; recommended: LhA (‘util/arc/LhA_e138.run’ on Aminet or

Fish disk 715) or Lx (‘util/arc/lx100.lha’ on Aminet), Unix version available

          (‘misc/unix/lha-1.00.tar.Z’)

.dms Disks compressed using DMS (‘util/arc/dms111.sfx’ on Aminet or Fish disk 406)

.zom Disks compressed using Zoom (‘util/arc/Zoom_5.4.lha’ on Aminet, Fish disk 682); an older version which you probably need for uncompressing PasTeX is found on Fish disk 459.

.zoo Compressed archive; recommended: Zoo (‘util/arc/zpp2-10.lzh’ on Aminet or Fish disk 527)

.Z

.z

.gz Compressed files; recommended gzip (‘util/arc/gzip-1.2.3bin.lha’ on Aminet), note that this are Unix files in most cases

.tar Acrhive; recommended: tar (‘util/arc/tar.lha’ or ‘util/arc/gtar10.lha’ on Aminet or Fish disk 445), note that tar is a Unix archiver and you often find soething

          like .tar.Z.

.arj Compressed archive; recommended unarj (‘util/arc/unarj-0.5.lha’ on Aminet)

.zip Compressed archive; recommended UnZip (‘util/arc/unzip-5.1.lha’ on Aminet), note that this are MS-Dos archives in most cases

7.3 Is there a Stacker-like utility to pack my hard drive?

XFH does a good job. It operates as a handler and uses the XPK-libraries, so you have different compression modes (NUKE is a good choice) and possibly even more in the future. The only disadvantage is, that the size of a file is limited by RAM: Don’t use it with less than 2MB of RAM.

Sources: Aminet (‘util/pack/XFH134.lha’ and ‘util/pack/xpk25usr.lha’), Fish disk 754.

Another possibility is EPU. It’s shareware and should offer the same as XFH, but without limiting the file size. Sources: Aminet (‘util/pack/epu14.lha’), Fish disk 858.

7.4 Where do I get Fish disk xxx?

Those FTP servers have that much space (or a CD-Rom) to have all fish disks available online:

     grind.isca.uiowa.edu    (Directory `/amiga/fx/fxxx')
     ftp.dfv.rwth-aachen.de  (Directory `/pub/amiga/fish')
     ftp.uni-erlangen.de     (Directory `/pub/amiga/pd')
     ftp.hawaii.edu          (Directory `/pub/amiga/fish')

Note that the CD-Rom’s are not always mounted. See Section 8.2 [FTP], page 19.

Another possibility would be to ask your local dealer. :-)

8 Where and how do I get Software?

Three questions arise in this context: Which programs can be found, where are they and how to get them and transfer them home?

8.1 Files and databases on freely distributable software

Of course you need to know where you find things. Many good choices are listed in this paper, as I hope. Other sources are:

AmigaSciSchool is a list of software and where you find it in Ascii format. It is posted monthly to the

          newsgroups comp.sys.amiga.applications, comp.unix.amiga, and news.answers.

Additionally you will find it on Aminet sites (‘text/doc/AmigaSciSchool-4.01’). It handles everything listed here and many more, for example GNU software, libraries (linked and shared), shells, Unix commands, educational and scientific software and much more.

FishCon are lists of the Fish disk contents. (‘fish/doc/fishcon-???.lzh’ on Aminet)

FishXref is a cross reference list of the Fish contents. (‘fish/doc/fishxref-???.lzh’ on

Aminet) FishXref and Fishcon are in Ascii format.

KingFisher A Fish disk database, (‘fish/doc/Kingfisher1_30.lha’, which is the program and ‘fish/doc/KFData850.lha’ ,which contains the data, on Aminet or Fish disk 863) allows search by name and context. See Section 8.4 [Fish], page 22.

8.2 Getting files from a FTP server

Things are easy for those who have access to the Internet and a program called FTP (File Transfer Program). Nearly all Unix computers have it, but not all of them allow the use of FTP.

FTP allows you to gain access to some other machines and store and/or retrieve files. Normally one needs an acoount on the remote machine to use it, bat a number of machines have a setup that allows everybody to log in as the user ftp or anonymous, so anybody may get files from them. The most important of this servers are the Aminet servers, which mirror each other and hence should have the same files. They are the best choice if you are looking for Amiga software. Aminet hosts are

     ftp.luth.se     (Sweden)
     litamiga.epfl.ch     (Switzerland)
     ftp.uni-kl.de     (Germany)

ftp.cs.tu-berlin.de     (Germany)
ftp.uni-erlangen.de     (Germany)
ftp.uni-paderborn.de     (Germany)
ftp.th-darmstadt.de     (Germany)
ftp.uni-oldenburg.de        (Germany)
wcarchive.cdrom.com     (USA)
ftp.etsu.edu     (USA)
ftp.wustl.edu     (USA)
merlin.etsu.edu      (USA)
oes.orst.edu     (USA)
src.doc.ic.ac.uk     (Great Britain)

All these mirrors have a directory ‘/pub/aminet’, where you will find much stuff. Please use a mirror close to you! Some other important hosts are

wuarchive.wustl.edu  (USA)
ftp.funet.fi  (Finland)
ftp.cso.uiuc.edu  (USA)
grind.isca.uiowa.edu  (USA)
ftp.dfv.rwth-aachen.de  (Germany)

Note that grind, aachen and erlangen have the full collection of Fish disks available! See Section 7.4 [Fish disk xxx], page 18.

To connect to a special host (ftp.uni-erlangen.de for example), you should type

ftp ftp.uni-erlangen.de

The host answers by requesting your login. You should type

ftp

No you are asked for a password. Please type your Email address here, if you have one. If not, use the password ftp.

Now you’re inside the host. There is a number of commands you may execute here. The most important are:

? Prints the help text of the FTP command. Additionally you may type ? command to get information on a special command.

bin Tells the FTP program that you whish to transfer binary files. It is always a good choice to type bin as the very first command! Files you load without the bin command can be corrupt.

get <file> Loads the given file from the host. On most Unix machines you can type something like ‘get file.txt -’ or ‘get file.txt |more’ to show a text on the screen. Note that there must be no blank between the | and the word more!

mget <pat> Loads the given files. pat may contain Unix style like wildcards. put <file> mput <pat>

Like get and mget, but transfer files from you to the remote host. This is in most cases not allowed, except for a special directory called ‘incoming’. You can place files here which you want to make public.

cd <dir> Like the usual cd command. The commands get, mget, put, mput, dir and ls refer to the current working directory.

dir [<dir>]

ls [<dir>] Like ‘list’ and ‘dir’ on the Amiga. Note that the FTP-dir corresponds to the Amigalist!

bye Leaves the FTP program.

When you have used FTP for the first times you will notice, that you always begin with executing the same steps:

1. Type the login (ftp in most cases)

2. Type the password (your mail address in most cases)

3. Enter the bin command

4. Change the current working directory (‘/pub/aminet’ for example)

This may get executed automatically. What you need is a file called ‘.netrc’ in your home directory. Note that it needs to be protected against others! The FTP program doesn’t use it, if it can be read by anything else than you. (Protection is set using the command ‘chmod go-rwx .netrc’.) The .netrc file contains some entries for your most favourite FTP sites, each separated by empy lines. A typicel entry may look like this:

     machine ftp.uni-erlangen.de
     login ftp
     password <your mail address> or <ftp>
     macdef init
         bin
         cd pub/aminet

Note that on some machines it is possible to use the machine name ‘default’ which meets all machines not listed in .netrc.

8.3 Getting files from a Mail server

Another way to get files is to use a mail server. This assumes that you can send mail to Internet addresses and get mail from Internet hosts. It works by sending a mail to the server specifying some commands, for example send commands for the files you want. Important mail servers are

     ftpmail@decwrl.dec.com
     mailserver@nic.funet.fi
     ftp-mailer@ftp.informatik.tu-muenchen.de
     mrcserv@janus.mtroyal.ab.ca
     mail-server@ftp.cs.tu-berlin.de
     mail-server@rtfm.mit.edu

The most important commands are:

Help Tells the server that you wish to get an Ascii file containing detailed explanation how

to use the server.

Limit <number> Specifys that you wish to get not more than <number> Kbytes per mail. Larger files get splitted into small pieces of at most <number> Kbytes which are sent as separate mails each. Note that the mails may get larger because of overhead.

Cwd <dir> Sets the current working directory to <dir>. This directory is used by the commands send and dir.

Index will return a list of files and/or directories that the server offers. Note that this may

be very large!

Index <item>

returns a list of files containing <item> in their names.

Dir [<dir>]

returns a list of the files and directories in the given directory.

Send <file1> <file2> … <fileN> Tells the server to send the given files to you.

Begin Tells the server to ignore all lines above this command.

End Like Begin, but specifies to ignore the lines below. (A signature for example!)

A typical session would be to send the following mail to the mail server:

     BEGIN
     CD /pub/aminet/util/arc
     SEND LhA_e138.run
     END

8.4 The Fish disks

A very good source are the Fish disks. One distinguishes between the floppy disks and the CD-Roms.

8.4.1 The Amiga Library disks

Fred Fish has started in the middle eighties to collect freely distributable software on floppy disks. There are more than 900 disks for now and very much good stuff on it. Most Amiga dealers sell them and most magazines contain addresses of people mailing them to you for about 3$ per disk or less.

Fred Fish has announced to terminate offering software on floppy disks with number 1000. Instead he offers CD-Roms. See Section 8.4.2 [Fish CD], page 23.

There are some things which can be found on the Fish disks, but not on Aminet. However, you can get them with FTP. See Section 7.4 [Fish disk xxx], page 18.

8.4.2 The Fresh Fish CD-Roms

Fred Fish is going on to offer freely distributable software. But now he collects it on CD-Rom’s. He will release two different kinds of CD’s:

1. Monthly released disks are divided into roughly three sections:

1. New material, which includes the material from the new unreleased floppy disks as well as material which does not appear in the floppy distribution, about 84Mb on the first disk.

2. Useful utilities that can be used directly off the CD-ROM if desired, thus freeing up the corresponding amount of hard disk space (GNU Emacs, Gnu C, GNU C++, Amiga E, PasTeX, AmigaGuide, Installer, 2.0 and 3.0-Includes, different archivers, tape drivers, the AmiCDROM filesystem and many other GNU and BSD tools, …), about 150Mb on the first disk.

3. Older material from previous released floppy disks or CD-ROM’s, about 404 Mb on the first disk. (Fish disk 600-910)

2. Disks containing the latest software as well as recent software in packed format only. (These are intended to be used for example in BBS’s.)

I recommend especially the first kind of Fish CD’s. They cost about 30$ plus 3$ for shipping (5$ outside USA/Canada/Mexico) and are available from

     Amiga Library Services
     610 N. Alma School Road, Suite 18
     Chandler, AZ 85224-3687
     U.S.A.
     Phone/FAZ: (602) 917-0917

8.5 How do I Read and write MS-Dos disks?

No problem for owners of Workbench 2.1 or higher: The program CrossDos is part of the Workbench. All you have to do is mounting the device ‘pc0:’ by putting it into the drawer ‘Devs:DOSDrivers’ or by double-clicking the icon in ‘Sys:Storage/DOSDrivers’. Ms-Dos disks in drive ‘df0:’ can now be handled in the usual manner replacing the word ‘df0:’ by ‘pc0:’. For example the directory can be shown with the command dir pc0:.

People still running Workbench 2.0 or lower need a program called MSH. You will find this on the Aminet (directory misc/emu) and on Fish disk 382. See Chapter 8 [Sources], page 19. After editing the file ‘devs:MountList’ as described in the documentation you have to say Mount msh: in the CLI and can now do the same as above replacing the word ‘pc0:’ with ‘msh:’.

8.6 How do I split large files?

There are some archives which are too large to fit on one disk. To transfer them on disks you need to split them into smaller pieces and transfer each part on a separate disk. I recommend Martin Schlodder’s Splitter. (Aminet, ‘util/misc/splitter_121.lha’. The archive contains binaries for MS-DOS and should be compilable without problems on any Unix system.

The Amiga-FAQ archive

The Amiga-FAQ is available in different formats: Ascii format (which is posted to the nets) AmigaGuide format (which is the adequate format on the Amiga) and in dvi format (to be printed). Additionally there is some stuff, that might be useful or interesting, but could not be included into the Amiga-FAQ:

     txt/amiga.history On the Amiga's history
     txt/story.txt The Commodore story (or: the Tramiel story ;-)
     txt/amiga.newsgroups Overview on comp.sys.amiga.*
     txt/amiga.sites  List of FTP sites
     txt/AmigaOverview.tex A short overview on the Amiga-Soft- and Hardware
     txt/Hardware.tips For those people who can't live without solder
     src/JWSplit.c The source of a file splitter
     src/JWJoin.c The opponent to JWSplit
     src/addtoc.c Utility to add a toc to texinfo-created docs
     (this document uses it)

I decided to collect these in the Amiga-FAQ archive. It is called AmigaFAQxxxxxx.lha (where xxxxxx is the date of the last release) and can be found on Aminet, directory ‘text/docs’.

Contributions

This FAQ can neither get useful nor hit further development without your help. Suggestions, contributions, new answers, critics, anything is rather welcome.

Please note, that very major subjects are absolutely missing yet: Nothing about sound, nothing on graphic cards, no Animation. These are some of the Amiga’s best points! But I don’t know them …:-(

So grab your keyboard (Your pencil? Well, if there’s no other way…) and send mail to:

           Jochen Wiedmann
           Am Eisteich 9
     72555 Metzingen (Germany)
           Tel. 07123 / 14881
           Internet: wiedmann@mailserv.zdv.uni-tuebingen.de

Credits

My thanks go to:

Reinhard Spisser and Sebastiano Vigna

for the Amiga version of texinfo. This is written with it.

The Free Software Foundation

for the original version of texinfo and many other excellent programs.

Dylan McNamee for contributing the sections on Editors, Word Processors, DTP and Postscript and some wording fixes.

Joseph Luk

for help in the section on chunky/planar, double buffering and PAL/NTSC

Urban Dominik Mueller for the FAQ on FTP and Mail-servers.

Index

.

.arj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .dms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .gz. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .lha. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .lzh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .netrc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 .tar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .z. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .zom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 .zoo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

6

68EC020. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 68EC030. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A

Amiga Library disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Amiga-FAQ archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 AmigaBasic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 AmigaMail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 AmigaSciSchool. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Anonymous. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Assemblers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 AutoDocs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

C

C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 CATS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Chunky displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Commodore, Frankfurt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Commodore, West Chester . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Compilers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Contributions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 CrossDos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

D

DaggeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Desktop Publishing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Developer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Doublebuffering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

E

Editors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Emulators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 endings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Enforcer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Esc sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

F

file endings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Fish CD-Rom’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Fish disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 Fish floppy disks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 FishCon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 FishXref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Forth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Fortran . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 FPU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Fresh Fish CD-Rom’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 FTP servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

G

GfxBase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GigaMem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

H

HD compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Hirsch & Wolf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

I

Includes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

K

KingFisher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

L

Linux. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

M

Mail-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Memory, virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 MMU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Modula-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Monitors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 MS-Dos disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Msh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Multiscan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

N

NDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NDUK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NetBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 NTSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

O

Oberon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

P

packers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 packers on Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Page Layout Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 PAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Pascal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Planar displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Postscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Printer control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Prolog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

R

RKM’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Rom Kernel manuals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

S

Splitting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Stacker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

T

TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Text Editors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

U

Unix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Unix-LhA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

V

VGA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

W

Word Processors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Wysiwyg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

X

X11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 XFH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 XPK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

Table of Contents

1 Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 What are the 68EC020 and the 68EC030? . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 What’s a FPU? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 The Operating System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2.1 Can I use another Kickstart than the builtin? . . . . . . . . . . . . . . . . . . . . . . 2

3 Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3.1 What documentation do I need as an Amiga programmer? . . . . . . . . . 2 3.2 What is CATS? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3 Where do I get the Amiga includes? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.4 How do I become a developer?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.5 What compilers (assemblers) are there? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.6 Those never working Esc sequences! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.7 Is it possible to use AmigaBasic on the A1200? . . . . . . . . . . . . . . . . . . . . 8

4 Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 Text Editors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2 What word processors are there? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3 Desktop Publishing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4 What is TeX and where can I get it? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.5 Are there any Postscript interpreters? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5 How about Graphics?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5.1 What are chunky and planar displays?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.2 What is doublebuffering? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.3 What monitors will work on my Amiga 1200/4000? . . . . . . . . . . . . . . . 14 5.4 How do I switch between PAL and NTSC? . . . . . . . . . . . . . . . . . . . . . . . 15

6 Emulators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

6.1 Can I run Unix on my Amiga? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 6.2 Is it possible to use the Amiga as X11 terminal? . . . . . . . . . . . . . . . . . . 17

7 Miscellaneous. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

7.1 Is there any unix version of LhA?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.2 What are files ending with ...? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.3 Is there a Stacker-like utility to pack my hard drive? . . . . . . . . . . . . . . 18 7.4 Where do I get Fish disk xxx? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

8 Where and how do I get Software? . . . . . . . . . . . . . . . . 19

8.1 Files and databases on freely distributable software . . . . . . . . . . . . . . . 19 8.2 Getting files from a FTP server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.3 Getting files from a Mail server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.4 The Fish disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

8.4.1 The Amiga Library disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8.4.2 The Fresh Fish CD-Roms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

8.5 How do I Read and write MS-Dos disks? . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.6 How do I split large files? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

The Amiga-FAQ archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

Contributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26